Prev Next

SW&IT Meeting Minutes 2019-10-07

Attendees: meeting adjourned

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

News from management

  • Oct 29, UCAP email: job title, salary range, etc.

Previous Action Items

Open topics

  • do we have adsec/wfs spares?
  • new AO/4D reduction VM server, AO data plan
  • PEPSI cluster
  • catalogs.lbto.org
  • iLocater 10Gb, snmp
  • FosWIKI deaths - reinstall new OS?

Team reports

DE

  • Raritan KVM test unit arrived
  • VMWare has approved our waiver; quote RSN
  • Opened converstation with Netops, ticket open
  • Power cut postmortem document
  • PEPSI dataflow sketch
  • AO hardware move
  • Upcoming OSX upgrade removes 32-bit compat libs; AnyConnect <3.8 is affected. Upgrade our ASA to supply newest version?
  • FACSUM work with AC
  • ARGOS BCU connection fix

MB

  • OC, WPC, management
  • review DE's conceptual PEPSI sketch
  • follow up misc Facsum/Argos issues
  • ALTA iterations, bad rear wind sensing (occacionaly random TBC), agreed some ways forward to detect the condition
  • AO data - DM overview, needs close scrutiny (volunteers?)
  • DMS
    • m1surf ready (for the plot)
    • conceptual tcs-snoop dds capture, iterations with PK
      • down the line, single place for streaming & debug (1 night) purpose - looking bright
  • safety trainings
next
  • DMS
    • live streams, alta, m1surf, tcs-snoop
    • learn about redis
  • catch-up with SH

CH

Last week:
  • CHAOS:
    • Emails to LINC-NIRVANA and LUCI PIs requesting info on alerting mechanisms. Requests forwarded to s/w leads.
    • Discussed option to adapt Zabbix as the CHAOS framework. MB asked for a paper showing what features we can leverage.
  • DMS: PG provided list of Monnit TM display mods he would like. Addressed by MB.
  • SupportScripts:
    • Refactored RB_Science.py, using pyds9 in place of subprocess.Popen.

This week:
  • CHAOS: Finish Zabbix paper.
  • SupportScripts: Package RB_Science as an RPM.

YZ

  • AO-Cravat:
    • Investigate plotly.js plotting for AO-Cravat data (testing and exploring several choices).
    • Started to redesign the whole web-interface.
    • Cleaning the AO-Cravat database a bit.
  • LBC/MODS:
    • Built the entire LBC telemetry database.
    • Reworking the LBC/MODS ingestion program so that they are robust (will not miss any telemetry data regardless of when they are run).
    • Need to finish the web-dms setups for LBC (involves new description for vectors)
    • Working on a scheme to fix the LBC non-sidereal guiding problem. Discussed with Olga. We have two schemes. I will investigate them and discuss with Olga again on how to proceed.
    • May spend some time on IT#7855 (MODS GUI problem). But may reassign it to Rick since I didn't find problems previously.

AC

  • IT7831 - removed -f full screen option to fix issue. Closed.
  • IT7882- updated spec file to include fitstopnm as dependency. Installed on NUCs. Need to test by taking daytime guider and wfs images. From GCS gui?
  • IT7876 - OAC suggests maybe perhaps socket flush or UMAC timeout?
  • IT7883 - OAC log indicates possible "-12v loss"
  • ARGOS
    • Updated SX BCU network configuration
    • Troubleshot and resolved connection to SX LAN controller issue
next
  • Vacation Monday
  • IT7876
  • IT7883

IS

Last week:
  • PIT
    • Created a number of xml2list scripts to create spreadsheets for partners and management.
    • Work on csv file for the archive people — two versions — one with all the proposals from start of PIT and one for 2019B.
  • OT
    • Work on point release that includes new functionality to allow passing warning/info messages from the backend to the OT everytime a script is generated
    • LBC calibration header
    • Provide a backward compatibility solution to the program to xml export/reimport MOS Mask bug introduced in OT 2020B

This week:
  • OT
    • LUCI ARC off lamps
    • FITS loading in PE issues
    • Tooltips for hi/low value cutoffs still not correct in sequence.
    • Look at DEC column in JSKY catalogs. It displays the wrong column.

  • LUCI
    • add to git /usr/local/bin/luci_geirs# where # is 1 or 2.
    • look at RTD MOS tab error calculation
Topic revision: r10 - 07 Oct 2019, MatthieuBec
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback